### v1

@@ Abstract
Regular black holes are often used to address singularities, but they typically involve a potentially unstable de Sitter core and an internal horizon that breaks predictability. Integrable singularities (IS) have recently gained attention because they avoid both issues and exhibit finite tidal forces, allowing nondestructive radial infall. First, we present a new BH solution sourced by a string fluid (FS) that exhibits an IS. Motivated by the divergence of the conserved energy in the cloud of strings (CS) model, we introduce an energy density profile based on the screening of the CS energy density within an FS framework, yielding a finite conserved energy. On the other hand, it has been proposed Ovalle (2024) that an interior region, rather than a pointlike mass, can generate a Schwarzschild BH exterior region. Secondly, motivated by the variety of BH solutions with singularities in the literature, we establish the conditions that an interior region with an IS must satisfy to represent a generic exterior BH solution, with Schwarzschild being only a particular case of the latter. We derive the junction conditions (JC) between the interior and exterior regions, showing that they lead to temperature continuity at the interface, while discontinuities in tangential pressure lead to phase transitions. We propose that the nature of the interior region is described by CS and FS, while the exterior corresponds to Reissner–Nordström.
@@ I Introduction
As an alternative to RBH solutions, Ref. Lukash and Strokov (2013) proposes a gravitational scenario in which a particle can reach the radial origin without being destroyed. Rather than a potentially unstable de Sitter core associated with an inner horizon, this model exhibits an integrable singularity at the origin. Integrable singularities are characterized by a divergence of the Ricci scalar while their volume integrals remain finite near the center. The elimination of the inner horizon thus prevents possible instabilities. Furthermore, recent studies in Ref. Estrada et al. (2024) show that a central integrable singularity allows for finite tidal forces, extendable radial geodesics, and a weak singularity, as formalized by Tipler’s extension theorem Nolan (2000). This ensures that an object approaching the central singularity would not undergo spaghettification. In addition, Ref. Casadio et al. (2023) offers a quantum perspective on the existence of such integrable singularities. For other recent applications, see Refs. Ovalle (2023); Estrada (2025).
@@
On the other hand, the Schwarzschild solution, which according to Birkhoff’s theorem Hawking and Ellis (2023) is the unique spherically symmetric vacuum solution of Einstein’s equations, possesses a strong central singularity. In this paragraph, we point out some ideas discussed in Ref. Ovalle (2024). The Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one. The interior region, defined by $r<h$, contains the central singularity shielded by the event horizon. In this domain, the metric function $f(r)$ becomes negative, causing the radial and temporal coordinates to interchange their roles and thereby revealing the dynamical character of the interior geometry. As a consequence, gravitational collapse toward the central singularity becomes unavoidable, in agreement with Penrose’s theorem Penrose (1965). Conversely, an observer in the exterior region has no causal access to the interior of the black hole: once the boundary at $r=h$ is crossed, it is impossible to return or communicate with the exterior. This Ref. Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$. In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere. In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core endowed with a potentially unstable inner horizon.
@@
It is well known that a wide class of black hole solutions studied in the literature are supported by matter sources that inevitably lead to a central singularity, for example BH cosmic-void density profiles Lustosa et al. (2025), BH dark-matter with dark matter energy profile Xu et al. (2018), and hairy BHs Ovalle et al. (2021). Electrically charged BH, exemplified by the Reissner–Nordström solution, also arise from electromagnetic sources in the energy–momentum tensor and are characterized by the presence of an inner horizon. Motivated by the above considerations, and in analogy with Ref. Ovalle (2024), we address as a second problem the existence of an internal geometry inside the event horizon of a generic black hole spacetime, keeping the external geometry completely general, with the Schwarzschild solution appearing only as a particular case. We focus on scenarios in which the internal structure is endowed with an integrable singularity, in order to avoid the emergence of a potentially unstable de Sitter core and the presence of an inner horizon, since predictability breaks down after crossing the inner horizon. We investigate the possible nature of matter sources capable of supporting such an internal geometry with an integrable singularity, specifically those associated with clouds of strings and with the new string fluid introduced in this work. To do this, we define the conditions that the internal geometry must satisfy in order to represent an integrable singularity, as well as the junction conditions between generic internal and external black hole geometries. Furthermore, we examine the connection between the Israel–Darmois Israel (1966) junction conditions and black hole thermodynamic properties, and explore their relation to the existence of phase transitions. As a test case, we study the Reissner–Nordström solution as an internal geometry, considering clouds of strings and the new string fluid as the first and second examples, respectively.

### v2

@@ Abstract
Regular black holes are often used to address singularities, but they typically involve a potentially unstable de Sitter core and an internal horizon that breaks predictability. Integrable singularities (IS) have recently gained attention because they avoid both issues and exhibit finite tidal forces, allowing nondestructive radial infall. First, we present a new BH solution sourced by a string fluid (FS) that exhibits an IS. Motivated by the divergence of the conserved energy in the cloud of strings (CS) model, we introduce an energy density profile based on the screening of the CS energy density within an FS framework, thereby yielding finite conserved energy. On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior region has recently gained attention Ovalle (2024); Maeda (2025). This is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon. Secondly, motivated by the variety of singular BH solutions in the literature, we establish the conditions that an interior region with an IS must satisfy in order to represent the interior of a generic exterior BH solution, with Schwarzschild being only a particular case of the latter. We derive the junction conditions (JC) between the interior and exterior regions, showing that they lead to temperature continuity at the interface, while discontinuities in tangential pressure lead to phase transitions. We propose that the nature of the interior region is described by CS and FS, while the exterior corresponds to Reissner–Nordström.
@@ I Introduction
As an alternative to RBH solutions, Ref. Lukash and Strokov (2013) proposes a gravitational scenario in which a particle can reach the radial origin without being destroyed. Rather than a potentially unstable de Sitter core associated with an inner horizon, this model exhibits an integrable singularity at the origin. Integrable singularities are characterized by a divergence of the Ricci scalar while their volume integrals remain finite near the center. The elimination of the inner horizon thus prevents possible instabilities. Furthermore, recent studies in Ref. Estrada et al. (2024) show that a central integrable singularity allows for finite tidal forces, extendable radial geodesics, and a weak singularity, as formalized by Tipler’s extension theorem Nolan (2000). This ensures that an object approaching the central singularity would not undergo spaghettification. In addition, Ref. Casadio et al. (2023) offers a quantum perspective on the existence of such integrable singularities. For other recent applications, see Refs. Ovalle (2023); Estrada (2025); Arrechea et al. (2025).
@@
On the other hand, the Schwarzschild solution, which according to Birkhoff’s theorem Hawking and Ellis (2023) is the unique spherically symmetric vacuum solution of Einstein’s equations, possesses a strong central singularity. In this paragraph, we point out some ideas discussed in Ref. Ovalle (2024). The Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one. The interior region, defined by $r<h$, contains the central singularity shielded by the event horizon. In this domain, the metric function $f(r)$ becomes negative, causing the radial and temporal coordinates to interchange their roles and thereby revealing the dynamical character of the interior geometry. As a consequence, gravitational collapse toward the central singularity becomes unavoidable, in agreement with Penrose’s theorem Penrose (1965). Conversely, an observer in the exterior region has no causal access to the interior of the black hole: once the boundary at $r=h$ is crossed, it is impossible to return or communicate with the exterior. In this line of research, remarkably, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior has recently gained attention Ovalle (2024); Maeda (2025). This configuration is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon, without introducing a lightlike thin shell. Reference Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$. In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere. In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core endowed with a potentially unstable inner horizon. Reference Maeda (2025) constructs a so-called fake Schwarzschild solution, observationally indistinguishable from the Schwarzschild black hole, by matching a dynamical spherically symmetric interior containing matter to the Schwarzschild exterior at the event horizon without introducing a lightlike thin shell. It is shown that this configuration can be obtained from a perfect-fluid solution satisfying a linear equation of state of the type considered in Ref. Maeda (2023), analytically continued inside the horizon.
@@
It is well known that a wide class of black hole solutions studied in the literature are supported by matter sources that inevitably lead to a central singularity, just to mention a few examples: black holes with cosmic-void density profiles Lustosa et al. (2025), black holes surrounded by dark matter with dark energy profiles Xu et al. (2018), hairy black holes Ovalle et al. (2021), and quintessential black holes Kiselev (2003); Estrada et al. (2025). Furthermore, electrically charged BH, exemplified by the Reissner–Nordström solution, also arise from electromagnetic sources in the energy–momentum tensor and are characterized by the presence of an inner horizon. Motivated by the above considerations, and in analogy with Ref. Ovalle (2024), we address as a second problem the existence of an internal geometry inside the event horizon of a generic black hole spacetime, keeping the external geometry completely general, with the Schwarzschild solution appearing only as a particular case. We focus on scenarios in which the internal structure is endowed with an integrable singularity, in order to avoid the emergence of a potentially unstable de Sitter core and the presence of an inner horizon, since predictability breaks down after crossing the inner horizon. We investigate the possible nature of matter sources capable of supporting such an internal geometry with an integrable singularity, specifically those associated with clouds of strings and with the new string fluid introduced in this work. To do this, we define the conditions that the internal geometry must satisfy in order to represent an integrable singularity, as well as the junction conditions between generic internal and external black hole geometries. Furthermore, we examine the connection between the Israel–Darmois Israel (1966) junction conditions and black hole thermodynamic properties, and explore their relation to the existence of phase transitions. As a test case, we study the Reissner–Nordström solution as an internal geometry, considering clouds of strings and the new string fluid as the first and second examples, respectively.